Solve: 2|2x -1| + 5 = 15
Answer by algebrahouse.com(1659)

2|2x -1| + 5 = 15
2|2x - 1| = 10 {subtracted 5 from both sides}
|2x - 1| = 5 {divided both sides by 2}
2x - 1 = 5 or 2x - 1 = - 5 {set that inside the absolute value equal to ±5}
2x = 6 or 2x = -4 {added 1 to both equations}
x = 3 or x = -2 {divided both equations by 2}
www.algebrahouse.com